The table below shows the maximum amount of civilian goods that an economy can produce for each level of military goods produced.

Point

Military Goods
(thousands of tonnes)

Civilian
(thousands of tonnes)

A

0

60

B

2

55

C

4

48

D

8

39

E

10

28

F

12

15

G

14

0

  1. Graph the PPC for this economy.
  2. If the economy is producing at point D, what is the opportunity cost of producing an additional 2 thousand tonnes of military goods?
  3. If the economy is producing at point E, what is the opportunity cost of producing an additional 2 thousand tonnes of military goods?
  4. Does the PPC satisfy the principle of increasing opportunity cost?

Question 2

Adam is in the middle of building a new concert venue in LA. The land and the initial construction materials have already been purchased and the framing has already been completed at a cost of $2 million. It will cost another $4 million to complete the project. Unfortunately, a new study suggests that demand for concert tickets is down and that competition from the Capitol Centre is expected to increase. Adam argues that he’s already spent $2 million and might as well complete the project. Is Adam mistaken? If so, how would you categorize any fault in his thinking?
Answer #1

PLEASE NOTE THAT I HAVE SOLVED FIRST 4 SUB PARTS OF THE FIRST QUESTION

(a)

(b) Consider that the economy is producing at D. To produce 2 thousands tonnes of military goods (i.e., to move to point E) , it has to give up 11 (i.e., 39-28) thousand tonnes of civilian goods. Hence, the opportunity cost of producing 1 thousand tonnes of military goods = 11/2 = 5.5 million tonnes of civilian goods.

(c) Now consider that the economy is producing at E. To produce 2 thousands tonnes of military goods (i.e., to move to point F) , it has to give up 13 (i.e., 28-15) thousand tonnes of civilian goods. Hence, the opportunity cost of producing 1 thousand tonnes of military goods = 13/2 = 6.5 million tonnes of civilian goods.

(c) The PPC does not satisfy the principle of increasing opportunity cost across all points. For example, if the economy is at B and it wants to move to point C, the opportunity cost of producing 1 thousand tonnes of military goods = (55-48)/(4-2) = 3.5. Further, when it wants to move from point C to D, the opportunity cost of producing 1 thousand tonnes of military goods = (48-39)/(8-4) = 2.5. Hence, from moving from C to D, opportunity cost declines. The tables below shows the opportunity cost of producing 1 thousand tonnes of military goods at all points. Except for the case of moving from C to D, the opportunity cost is increasing. The left side table shows the values of opportunity cost where as the right side table shows the formula view of the calculation of opportunity cost.

To satisfy the principle of increasing opportunity cost, the PPC has to be concave throughout. As can be seen from the diagram above (a), the concavity breaks in the region C to D.